OpenStudy (anonymous):

1. If y varies inversely as x, and y = 4 as x = 8, find y for the x-value of 2. 2.If y varies inversely as x, and y = 3 as x = 8, find y for the x-value of 6.

OpenStudy (johnweldon1993):

\[\large y = \frac{k}{x}\] Will be the formula we will utilize here..varying inversely

OpenStudy (johnweldon1993):

When y = 4....x = 8...so \[\large 4 = \frac{k}{8}\] Solving for 'k' we get \[\large k =32\]

OpenStudy (johnweldon1993):

If that is true...our formula for this problem will be \[\large y = \frac{32}{x}\] So now...when x = 2...what does 'y' equal?

OpenStudy (ibbutibbu.):

ok.. this is proportions. so... 1: The answer is Y=0.5 2: The answer is Y=2.25

OpenStudy (johnweldon1993):

We would plug in 2 for 'x' and solve for 'y' getting a result of \[\large y = \frac{32}{2} = 16\]